equation of a line
(-2,-8) and (-7,-5)
I got y=3/5x+-2/5
Is this right?

Answer by Alan3354(69443)

equation of a line
(-2,-8) and (-7,-5)
I got y=3/5x+-2/5
Is this right?
=======================
If you mean the line thru the 2 points,
and if it's y=(3/5)x - 2/5
------
Sub for x & y for the 2 points and see.
----------------
y = (3/5)x - 2/5
-8 =? (3/5)*-2 - 2/5
-8 =? -6/5 - 2/5
-8 =? -8/5
Doesn't work.
No reason to check the other point.